1. "Motion is relative” Justify this statement with suitable examples (2 eg)

2. An athlete starts running from the centre of a circular track of radius 7m along the

radius. On reaching the track he runs clockwise along the track upto a point diagonally opposite to his starting point. Calculate the distance travelled and displacement.

3. A ball is thrown vertically upwards to a height h and caught back.What is the

distance travelled and displacement?

4. Which is faster- a car running at 72 kmph or a ball rolling at 25 m/s?


1
Expert's answer
2021-06-29T09:59:42-0400

1) Yes, motion is always a relative term. A body can be said to be in motion , if it changes its position with respect to another body. Now say, if two bodies with moving with same velocity, in direction parallel to each other then for an observer in any of the two bodies , the other body will appear to be stationary.

2)


"d=\pi r=7\pi=22\ m\\s=2r=2(7)=14\ m"

3)


"d=h+h=2h\\s=h-h=0\\"

4)


"72\ kmph=\frac{72}{3.6}=20\frac{m}{s}<25\frac{m}{s}"